### v1

# LionHeart: A Layer-based Mapping Framework for Heterogeneous Systems with Analog In-Memory Computing Tiles

Abstract
When arranged in a crossbar configuration, resistive memory devices can be used to execute Matrix-Vector Multiplications, the most dominant operation of many Machine Learning (ML) algorithms, in constant time complexity. Nonetheless, when performing computations in the analog domain, novel challenges are introduced in terms of arithmetic precision and stochasticity, due to non-ideal circuit and device behaviour. Moreover, these non-idealities have a temporal dimension, resulting in a degrading application accuracy over time. Facing these challenges, we propose a novel framework, named LionHeart, to obtain hybrid analog-digital mappings to execute Deep Learning (DL) inference workloads using heterogeneous accelerators. The accuracy-constrained mappings derived by LionHeart showcase, across different Deep Neural Networks (DNNs) and datasets, high accuracy and potential for speedup. The results of the full system simulations highlight run-time reductions and energy efficiency gains that exceed 6$\times$, with a user-defined accuracy threshold with respect to a fully digital floating point implementation.
Index Terms:
AIMC, DNNs, Heterogeneous Systems, Mapping, System Simulation
I Introduction
Deep Learning-based solutions have been applied in various industrial and scientific applications, including computer vision, natural language processing, and speech recognition. Nowadays, State-Of-The-Art (SOTA) DL models require significant memory to store their parameters (i.e., weights and biases), and billions of Multiply-Accumulate (MAC) operations per inference. These extreme computing and memory requirements pose a challenge for deploying DL workloads on edge devices, whose form factor and energy budget require limited computing capabilities and memory size. However, the execution of DL workloads at the edge represents a very attractive prospect, as it would allow the processing of data samples close to where it is collected, resulting in several benefits, including reduced latency, a higher degree of security and privacy, improved bandwidth efficiency, as well as increased resiliency. Therefore, many research works have proposed strategies to optimize DL workloads for edge scenarios.
Figure 1: LionHeart considers the mapping of FC and CONV DNN layers to heterogeneous PUs featuring AIMC tiles and general-purpose cores to maximize performance while adhering to a user-specified accuracy constraint.
ML algorithms, such as DNNs and transformers, are structured as regular sequences of operations. For the most part, they implement linear algebra computing kernels, which can be effectively parallelized. Hence, a number of customized hardware architectures dedicated to accelerating ML workloads have been proposed in recent years. They exploit reprogrammable logic or fixed-function implementations, such as Tensor Processing Units. A disruptive approach in this space is In-Memory Computing (IMC), for which storage and computation is performed at the same physical location, avoiding the well-known memory wall problem.
Different IMC approaches have been proposed in the literature, operating at varying levels of the memory hierarchy and leveraging different technologies. Among them, Analog In-Memory Computing (AIMC) is emerging as a particularly appealing alternative. A key drawback hampering the widespread adoption of AIMC is its sensitivity to device and circuit-related variations such as temporal conductance variations, device stochasticity, and circuit non-idealities, which can adversely affect accuracy [1]. Many ongoing efforts investigate these effects and propose strategies to mitigate them, including Hardware-Aware (HWA) training [2] and model co-optimization [3].
Addressing the accuracy challenge from a system perspective, we present herein a novel framework that enables the effective exploration of hybrid digital/analog mappings of DL models (Fig. 1), taking into consideration the current limitations of AIMC which affect accuracy. Our framework effectively navigates the trade-off between run-time, performance, and accuracy degradation. Recognizing the exponential relation between the number of MVMs that can be executed in either the digital or analog domain, and the possible mapping solutions, we herein introduce an accuracy-driven heuristic that, by considering the largest MVM operations first (in terms of MAC operations), reduces the exploration problem to that of linear complexity. The main contributions presented in this work are summarized as follows:
  • We present a novel accuracy-driven training framework, able to explore the space of hybrid digital/analog implementations of DNNs. The framework allocates Fully Connected (FC) and Convolutional (CONV) layers to digital or analog resources to minimize run-time while constraining the adverse effect of analog computations on accuracy. Additionally, it is hardware-agnostic, meaning it can be applied to AIMC crossbars adopting different device technologies;
  • We derive a set of optimized hybrid mappings ensuring user-defined accuracy levels for a number of DNNs trained for CIFAR-10 and CIFAR-100 image classification, showing that in all cases a sizable share of their workload can be executed in the analog domain, even for degradation thresholds as low as 0.5% with respect to floating-point precision implementations;
  • We investigate the speedup-accuracy tradeoff achieved by our mappings. To this end, we perform power and performance evaluations on different types of ML workloads using a cycle-accurate simulator based on gem5-X. The considered hybrid system employs a single-core in-order CPU embedding tightly-coupled AIMCs for in-memory acceleration;
  • We recognize and address the real-world system accuracy degradation caused by temporal variation of programmed AIMCs crossbars at a user-desired evaluation time, $t_{eval}$.
II Background and Related Work
II-A In-Memory Computing
Traditional von Neumann computing systems involve separate processing and memory units, which consume significant energy and introduce additional latency when data is moved between them. IMC is a non-von Neumann computational approach, in which computation and storage are performed at the same physical location. From a technology perspective, both SRAM- and DRAM-based IMC architectures have been proposed [1]. Computation in these architectures can be achieved by leveraging charge sharing when concurrently activating multiple memory rows. A promising alternative, which we focus on in this paper, is that of employing Non-Volatile Memory (NVM) devices, organized in a crossbar arrangement [1]. These devices, used to encode network weights in a differential mapping scheme, are employed at the junction of word- and bit-lines, and their resistances are modulated according to the target weight value at compile time.
In AIMC crossbars, inputs are converted to word line voltages using Digital-to-Analog Converters, and currents, representative of MVMs results, are collected across bit lines into Analog-to-Digital Converters. Operations on AIMC crossbars are performed in parallel, and hence, can be realized in constant time complexity. For instance, a 256$\times$256 Phase Change Memory (PCM) crossbar is shown to perform a total of 65,536 MAC operations in 130ns [4]. However, as computation is performed in the analog domain, outputs are affected by non-deterministic noise, deviating from expected results. This is due to a number of device and circuit non-idealities, including temporal and temperature-dependent conductance variations, device stochasticity, ADC mismatch, and IR drop [5].
Figure 2: (a) High-level overview of our proposed framework. (b) The validation accuracy at the desired evaluation time ($t_{eval}=$1d) and training loss of a candidate AIMC layer during analog HWA retraining. Training is considered converged when, for a user-defined convergence window, the training loss does not decrease. (c) High-level overview of analog HWA retraining.
II-B Hardware-Aware Training
Analog HWA offline training is a common approach to mitigate some of the AIMC device and circuit non-idealities, by making the model more robust for deployment. Additional operations, including weight-noise injection, are added during forward propagation passes. Mathematically, the generally simulated AIMC forward passes can be expressed as
$y_{i}=\alpha^{\textrm{out}}_{i}f_{\textrm{adc}}\big{(}\sum_{j}({w}_{ij}+\sigma_{\textrm{w}}\xi_{ij})(f_{\textrm{dac}}(x_{j})+\sigma_{\textrm{inp}}\xi_{j})+\sigma_{\textrm{out}}\xi_{i}\big{)}+\beta_{i},$ (1)
where $f_{\text{adc}}$ and $f_{\text{dac}}$ model the analog-to-digital and digital-to-analog processes, with dynamic scaling and range clipping, and $\xi$ denotes Gaussian noise. $\sigma$ indicates noise sampled from a zero-centered normal distribution. In addition to training noise-resilient weights, it is also possible to train other parameters, such as the input (DAC) bounds of each Word-Line (WL), using back-propagation. We employ the open-source AIHWKIT toolkit [5, 6] to expose these effect in system-level explorations in LionHeart, allowing both the realistic characterization of analog non-idealities and analog HWA retraining. While the toolkit, and by extension, the LionHeart framework can be used to model other device technologies, such as Resistive Random-Access Memory (RRAM), in this paper, we model PCM. Section IV-B3 details the specific assumptions made during analog HWA retraining.
II-C Related Work
Related works, which perform application mapping on AIMC- accelerated systems, consider resource limitations as their main (or only) constraint. PUMA [7] employs compiler analysis passes to define which parts of the application have the most potential for speedup, while AERO [8] formulates this same problem as a cost function minimization. Potential issues introduced by non-idealities and noisy environments, characterizing AIMC devices, are overlooked in the above-mentioned works. They are similarly disregarded in [9] and [10], where all FC and CONV layers are executed using AIMC tiles, without attempting to control induced accuracy degradation. Some consideration of the effect of analog computing on accuracy is included in [11], where the authors propose a hybrid solution where the first and last layers of DNNs are executed on digital resources, postulating that these two layers are more sensitive to noise. Two recent works present input channel-wise weight selection methods to map to heterogeneous analog-digital resources. The first, HybridAC [12], does so without employing any re-training strategy. The input is assumed to already be HWA trained, and some layers are confined to be implemented using digital resources. The second, ODiMO [13], relies on the unique approach of training a supernetwork. Analog computation is modelled using 2-bit quantization. However, both of these methods do not consider (i) the variability of analog tiles as well as (ii) the temporal evolution of analog weights.
The complexity of devising high-performance, but accuracy-constrained mappings, is compounded by the varying size and sensitivity towards noise of the different layers in a DNN structure. Moreover, the number of possible mappings makes exhaustive exploration impractical, even for simple cases, as discussed, in the context of hardware quantization, in [14]. For example, even when layer-wise mapping is considered, a relatively low-depth model like AlexNet, composed of just seven layers, would therefore present $2^{7}$ possible analog-digital combinations. These would all have to be trained and tested to arrive at an exact solution to the mapping problem. In contrast to the aforementioned works, we employ a heuristic-based layer-wise mapping strategy which scales linearly with respect the number of layers, as it only consider one target candidate layer at a time.
III LionHeart framework
[…]
TABLE I: The number of layers, weights, and the top-1 (baseline, floating-point) accuracy of each evaluated network.
Model Weights Top-1 Acc. (%)
(# Mappable Layers)
CIFAR-10
CIFAR-100
CIFAR-10
CIFAR-100
ResNet8 (10) 77,360 77,968 87.38 59.40
ResNet20 (20) 268,336 270,522 90.52 65.15
MobileNet (15) 3,195,328 3,206,282 90.15 63.05
VGG-16 (14) 14,715,584 14,761,664 92.79 69.51
AlexNet (7) 15,378,112 15,562,432 91.29 65.70
[…]
IV Evaluation
IV-A Experimental Setup
The CIFAR-10 and CIFAR-100 datasets were chosen for evaluation. We collect results using these datasets for the five DNN architectures targeting edge devices as listed in Table I, where the size of the last FC layer is modified according to the number of output classes (10 and 100, respectively). The number of weights of each DNN ranges from 77,968 to 15,237,698, and hence contains both under- and over-parameterized models for the tasks. Data is split into training, validation, and testing sets. The validation dataset was obtained by splitting the original training dataset into two separate datasets, comprising 90% of training images and the remaining 10% of validation images. The same randomly sampled validation set was used for all experiments.
Figure 3: (a-e) The MAC ratio (as a percentage) for all configurations between the number of MACs performed using AIMC and the total number of MACs. (f-j) The corresponding test set accuracies at $t=$1d.
IV-B Training and Evaluation Strategies
IV-B1 Initial hyper-parameter exploration:
Our framework assumes that the input DNN is pre-trained in floating-point precision. The selected models differ in depth, size, and complexity, and hence require different hyperparameters to achieve optimal training performance. We empirically determined two sets of hyper-parameters, pertaining to the analog and digital domain, respectively. Then, these were applied to all networks. For analog training, a learning rate of 0.024 and momentum of 0.775 was used. In terms of digital training, we employed a learning rate of 0.057 and a momentum of 0.867. A convergence window of 5 epochs was used for retraining. Further details on the adopted experimental setups in the two domains are provided in the following.
IV-B2 Digital training:
Networks were trained using floating-point precision until the training loss did not decrease for five consecutive epochs. The batch size was set to 256. Stochastic Gradient Descent (SGD) with momentum was used, with cross-entropy loss. A cosine-annealing learning rate scheduler was used.
IV-B3 Analog HWA re-training:
To perform analog HWA re-training, the weights of the trained digital networks were employed as a starting point. The IBM AIHWKIT was used to inject noise during forward propagations. We refer the reader to [6] for a comprehensive tutorial on HWA training using IBM AIHWKIT. The InferenceRPUConfig() preset was used with the PCMLikeNoiseModel phenomenological inference model. The following additional modifications were made to the simulation configuration: (i) biases are assumed to be digital (i.e., they are not encoded on the last column of AIMC tiles). (ii) Channel- (i.e., column-) wise weight scaling, which has a negligible performance impact, is performed, where weights are mapped to a per-channel conductance range during learning after each mini-batch. (iii) The size of each AIMC tile is assumed to be 256x256. (iv) Layers that span across multiple AIMC tiles are assumed to be split evenly. (v) During training, multiplicative Gaussian noise is applied to unit weight values, with a standard deviation of $0.08$, extracted from hardware measurements provided by the authors of [4].
IV-B4 Accuracy evaluation:
As AIMC hardware is inherently stochastic, multiple evaluation instances are required to determine representative performance metrics. Consequently, accuracy was measured 20 times and averaged across experiments to obtain the values reported in Section IV-D. We considered five user-defined accuracy thresholds of [0.5%, 1.0%, 3.0%, 5.0%, 10.0%]. Moreover, analog NVM devices such as PCM are susceptible to temporal conductance drift. Since the behavior of AIMC hardware can evolve over time, we explore in Section IV-F accuracy metrics for AIMC accelerators programmed considering drift levels predicted at [one second (i.e., $t=1s$), one minute (i.e., $t=60s$), one hour (i.e., $t=3600s$), one day (i.e., $t=86400s$)]. These are then evaluated at different points in time, to assess the effect of a mismatch between predicted and actual drift.
IV-B5 Full System Evaluation:
The target simulated system for this work is a high-performance edge processor with an in-order ARM CPU core running at a frequency of 2.3GHz, with a 64KB each L1 data and instruction caches, 1MB of L2 cache, and 8GB of 2400MHz DDR4 RAM. The software stack for the target DNNs includes the DNNs accelerated via the Eigen C++ library, running in user-space atop the Linux 5.4 kernel, and an Ubuntu 16.04 LTS disk image. Each implementation resulting from a LionHeart mapping is run for 10 inferences to reduce simulation noise. We note that measured deviations among different runs of the same mappings never exceeded 4% of the run-time.
IV-C Baselines
LionHeart (in all plots, referred to as LH) is evaluated by comparing the achieved accuracy and performance with respect to three baselines: (i) Fully-digital. The floating-point DNN model input of our methodology, where none of its layers are accelerated using AIMC. It achieves the highest accuracy, but presents the longest run-time. (ii) Fully-analog. A DNN implementation with all FC and CONV layers executed using AIMC. This baseline represents the opposite of the fully-digital: it maximizes performance, while potentially resulting in very large accuracy degradation. (iii) First-Last Mapping Strategy (FLMS). The SOTA hybrid mapping strategy [11]. In these mappings, only the first and the last layers are implemented in digital PUs, while all others are AIMC-accelerated.
IV-D Accuracy Evaluation
For five different drop thresholds and baseline configurations, we determined the ratio between MACs performed using AIMC and the total number of MACs (Fig. 3). As can be observed, for all networks, the All Digital configuration has the highest test set accuracy and the All Analog configuration has the lowest test set accuracy. The behavior of the FLMS configuration, however, is unpredictable. The proposed LionHeart configurations exhibit a gradual transition from the largest to the smallest test set accuracy and a corresponding decrease in the digital/analog MAC ratio between all configurations. It hence enables the exploration of the performance/accuracy space exposed by AIMC acceleration, while controlling accuracy to a user-defined constraint.
For 62% of configurations, the last FC layer is retained in analog. However, for only 18%, the first CONV layer is retained. This means that the first layer is usually the most sensitive. Intermediary layers are observed to be more sensitive than the last layer. For intermediary layers, no strong trend or indication is observed as to whether a layer will remain analog, other than that all FC ones are more likely to be retained as digital layers. Hence, it is likely, for CONV layers, that whether or not a layer will be retained in analog is largely dependent on the trained weights and not on other layers’ characteristics. Our analyses demonstrate that properly mapping layers to digital or analog resources is non-trivial, showcasing the benefit of our framework in comparison with more naïve strategies such as FLMS.
IV-E Performance Evaluation
TABLE II: Architectural parameters for full system simulation.
CPU Model 1x In-Order @ 2.3GHz
ISA ARMv8
Supply Voltage VDD 1.3V
L1 Data/Instruction Cache 64kB
L2 Cache 1MB
Memory 8GB DDR4 @ 2400MHz
AIMC Tile Latency 100ns
AIMC Tile Data Transfer Bandwidth 4GB/s
AIMC Tile Energy Efficiency 20 TOp/s/W
Figure 4: (a) Speedup and (b) energy efficiency gain resulting from different application mapping strategies, when performing an inference on benchmark DNNs. Data is normalized with respect to a fully digital execution with no AIMC acceleration.
In LionHeart, the workload associated with DL models and their layers is assumed to be architecture-agnostic, i.e., the ratio of analog MAC to total MAC operations is used as a proxy for speedup and energy efficiency. To quantify how strongly the MAC ratio correlates to these metrics, we evaluate them for a simulated system featuring a single-core in-order CPU clocked at 2.3GHz and a two-level cache hierarchy. The architecture embeds tightly-coupled AIMCs, governed by dedicated extensions to the ARM instruction set supported by the CPU. It is modeled using the ALPINE framework [15], allowing the full system simulation of AIMC-accelerated systems. Table II lists the adopted architectural parameters, which mirror the ones considered in [15] for the single-core, higher-performance scenario.
Due to the L1 and L2 cache sizes (rendering AIMC-acceleration of smaller networks ineffective), results were gathered on the three largest DNNs among the ones presented in Section IV-D, considering the more challenging CIFAR-100 dataset. We simulated the execution of an entire inference, including computations not pertaining to FC nor CONV layers. We investigated heterogeneous mappings derived by LionHeart, with varying maximum accuracy degradation thresholds. For comparison, we also conducted experiments where all FC and CONV layers are completely executed on AIMCs, and mappings where, similar to [11], all layers except the first and last are hosted on analog crossbars (FLMS).
The complete simulation outcomes of the system, presented in Fig. 4, highlight that the achievable speed-ups are highly dependent on the DNN structure. In particular, even when all layers are computed in the analog domain, which is an upper bound for speedup and efficiency in our setting, only a speedup of 24% can be reached in the MobileNet benchmark. Such (comparatively) low run-time performance improvement is caused by the lower degree of data reuse of the depthwise separable convolutions employed in this network, which shift the balance between computation and memory access requirements toward the latter, limiting the benefit of accelerating MACs. Nonetheless, even in this case, LionHeart, with a limited maximum accuracy degradation of 5%, still captures almost all of the available speedup and energy efficiency gains.
Figure 5: The impact of temporal drift when varying $t_{eval}$ in [0, 86400] for the ResNet8 network evaluated on CIFAR-100.
AlexNet and VGG16, which do not employ separable convolutions, present instead more potential speedup (660% in the full-analog case for VGG16, 550% for AlexNet), due to the higher locality of the employed computations (which leads to less cache misses and memory-induced stalls). Even in those cases, LionHeart can harness a large part of the potential speedup if enough accuracy leeway is allowed. For a 5% accuracy degradation, a 425% speedup was measured in AlexNet and a 285% one in VGG16. AlexNet proved to be the most robust of the investigated networks. In fact, for a 10% accuracy degradation threshold, its performance exceeded that of the FLMS baseline, as all layers can be executed in AIMCs, further showcasing the benefit of adopting a flexible accuracy-driven approach. Finally, it can be noticed that system-level energy gains closely follow speedups. Indeed, AIMCs tiles are not a major contributor to overall system power (as opposed to processors and memories). Instead, they shorten run-time, ultimately lowering energy requirements. This is achieved both by allowing the execution of MVMs in constant time and by encoding DNN weights in the crossbar structure itself, hence easing the pressure on the memory hierarchy.
IV-F Temporal Drift and Evaluation Time Analysis
We study the impact of varying the evaluation time, $t_{eval}$, used during analog HWA retraining on the simulated inference time of the system. In Fig. 5, these metrics are reported for all configurations (except All Digital, which does not have a temporal dimension) using the ResNet8 network and CIFAR-100 dataset. For both sets of mappings, mostly unique intermediate layers are retained in analog. Due to the (relatively) small drop threshold, the first and last layers for each are in digital. It can be observed that the test accuracy decreases a function of the drift time. Moreover, as the proportion of analog MACs increases, the robustness to drift decreases. Most notably, for the All Analog configuration, independent of the evaluation time, the test accuracy decreases significantly (¿10%). For all other configurations, no significant degradation is observed until $t$¿1h. However, the temporal dependence of the test set accuracy is effected. Unsurprisingly, when a larger evaluation time is specified, the initial test set accuracy (at $t$=0) is also larger. While varied, for most configurations the robustness to drift is observed to improve as the evaluation time is increased. This is significant because, when our methodology is used, a hard constraint on the average behavior of the drop in accuracy can be enforced at a given drift (evaluation) time. Additionally, the test set accuracy for the FLMS configuration is consistently lower than for all LionHeart configurations.
V Conclusion
In this paper, we presented LionHeart, an accuracy-driven framework for mapping DL workloads on hybrid analog-digital systems. The effectiveness of our method was demonstrated by deploying five different DNNs trained for image classification to heterogeneous PUs. Both the MAC ratio between analog and digital operations and the effect of temporal drift were considered to enforce varying user-specified constraints. Finally, the performance evaluation was carried out using a simulated ARMv8-based processor with tightly coupled AIMC tiles.
VI Acknowledgements
This work was supported by the European Commission under H2020 grant WiPLASH (GA 863337), the EC H2020 FVLLMONTI project (GA No. 101016776), and by the ACCESS – AI Chip Center for Emerging Smart Systems, sponsored by InnoHK funding, Hong Kong SAR. We also thank the IBM Research Analog AI team for their feedback and acknowledge computational support from AiMOS, an AI supercomputer made available by the IBM Research AI Hardware Center and Rensselaer Polytechnic Institute’s Center for Computational Innovations (CCI).