### v3

# LionHeart: A Layer-based Mapping Framework for Heterogeneous Systems with Analog In-Memory Computing Tiles

Abstract
When arranged in a crossbar configuration, resistive memory devices can be used to execute Matrix-Vector Multiplications, the most dominant operation of many Machine Learning (ML) algorithms, in constant time complexity. Nonetheless, when performing computations in the analog domain, novel challenges are introduced in terms of arithmetic precision and stochasticity, due to non-ideal circuit and device behaviour. Moreover, these non-idealities have a temporal dimension, resulting in a degrading application accuracy over time. Facing these challenges, we propose a novel framework, named LionHeart, to obtain hybrid analog-digital mappings to execute Deep Learning (DL) inference workloads using heterogeneous accelerators. The accuracy-constrained mappings derived by LionHeart showcase, across different Convolutional Neural Networks (CNNs) and one transformer-based network, high accuracy and potential for speedup. The results of the full system simulations highlight run-time reductions and energy efficiency gains that exceed 6$\times$, with a user-defined accuracy threshold for a fully digital floating point implementation. LionHeart is open-sourced here: https://github.com/IBM/lionheart.
Index Terms:
AIMC, DNNs, Heterogeneous Systems, Mapping, System Simulation
I Introduction
Deep Learning (DL)-based solutions have been applied in various industrial and scientific applications, including computer vision, natural language processing, and speech recognition. Nowadays, State-Of-The-Art (SOTA) DL models require significant memory to store their parameters (i.e., weights and biases), and billions of Multiply-Accumulate (MAC) operations per inference. These extreme computing and memory requirements pose a challenge to deploy DL workloads on edge devices, whose form factor and energy budget require limited computing capabilities and memory size. However, the execution of DL workloads at the edge represents a very attractive prospect, as it would allow the processing of data samples close to where it is collected, resulting in several benefits, including reduced latency, a higher degree of security and privacy, improved bandwidth efficiency, as well as increased resiliency. To enable DL within the tiny resource envelopes of edge devices, workload optimization is key. To this end, many research works have proposed dedicated hardware solutions, customized for DL workloads [1, 2].
Figure 1: LionHeart heterogeneously maps layers of ML networks to digital or analog resources, maximizes performance by exploiting AIMC acceleration, while at the same time abiding to accuracy constraints.
Such accelerators leverage the structured organization of ML algorithms, such as Deep Neural Networks and transformers, which are composed of layers sequences that, for the most part, implement linear algebra computing kernels. In turn, these can be effectively supported by hardware architectures by exploring their parallelism, whether employing reprogrammable logic or fixed-function implementations, such as Tensor Processing Units [3]. A disruptive approach in this space is In-Memory Computing (IMC), for which storage and computation is performed at the same physical location, avoiding the well-known memory wall problem.
Different IMC approaches have been proposed in the literature, operating at varying levels of the memory hierarchy and leveraging different technologies. We summarize this landscape in Section II. Among them, Analog In-Memory Computing (AIMC) is emerging as a particularly appealing alternative. The AIMC computing paradigm is based on crossbar architectures, with devices having variable conductance connecting rows and columns. By leveraging Ohm’s law and Kirchhoff’s current law, AIMC accelerators can be used to perform matrix-vector operations between an input vector (encoded using Word-Line (WL) voltages) and a matrix (encoded using conductance) in linear time [4], i.e., the time required to write inputs and read-back outputs.
An important drawback hampering the widespread adoption of AIMC is its sensitivity to device and circuit-related variations such as temporal conductance variations, device stochasticity, and circuit non-idealities, which can adversely affect accuracy [2]. Many ongoing efforts investigate these effects and propose strategies to mitigate them, including Hardware-Aware (HWA) retraining [5] and model co-optimization [6]. However, none of these works adequately addresses the accuracy challenge from a system perspective, i.e. how AIMC acceleration can be leveraged to speed up ML execution while faithfully modeling and controlling the ensuing noise-induced accuracy degradation.
We herein aim at filling this gap. To this end, we present a novel framework that enables the effective exploration of hybrid digital/analog mappings of DL models (Fig. 1), taking into account the current limitations of AIMC which affect accuracy. Our framework effectively navigates the trade-off between run-time, performance, and accuracy degradation. Recognizing the exponential relation between the number of MVMs that can be executed in either the digital or analog domain, and the possible mapping solutions, we herein introduce an accuracy-driven heuristic that, by considering the largest MVM operations first (in terms of MAC operations), reduces the exploration problem to that of linear complexity. The main contributions presented in this work are summarized as follows:
  • We present a novel accuracy-driven training framework, able to explore the space of hybrid digital/analog implementations of DNNs. The framework allocates different layers to digital or analog resources to minimize run-time while constraining the adverse effect of analog computations on accuracy. Additionally, it is hardware-agnostic, meaning it can be applied to AIMC crossbars adopting different device technologies;
  • We derive a set of optimized hybrid mappings ensuring user-defined accuracy levels for a number of DNNs trained for CIFAR-10 and CIFAR-100 image classification, in addition to a transformer-based network for Stanford Question Answering Dataset (SQuAD) – a Natural Language Processing (NLP) task;
  • We showcase, that for all of these networks, a sizable share of their workload can be executed in the analog domain, even for degradation thresholds as low as 0.5% with respect to floating-point precision implementations;
  • We investigate the speedup-accuracy tradeoff achieved by our mappings. To this end, we perform power and performance evaluations on different types of ML workloads using a cycle-accurate simulator based on gem5-X [4];
  • We compare our optimal mappings to those obtained from prior work (DIANA [7] and Harmonica [8]), and evaluate the variability of the achieved mappings, in addition to their accuracy and MAC ratio;
  • For one network, we hardware-aware train all possible mappings and establish a upper-bound for comparison;
  • Finally, we recognize and address the real-world system accuracy degradation caused by temporal variation of programmed AIMCs crossbars at a user-desired evaluation time, $t_{eval}$.
II Background and Related Work
II-A In-Memory Computing
Traditional von Neumann computing systems involve separate processing and memory units, which consume significant energy and introduce additional latency when data is moved between them. IMC is a non-von Neumann computational approach, in which computation and storage are performed at the same physical location. Several different IMC architectures, which aim to blend computation and storage, have emerged in recent years [9]. Architectures can utilize IMC at different levels of the memory hierarchy, for example, when interfacing main memory [10], as part of smart caches [11], or as functional units that augment processor pipelines and register files [4]. While we consider this last scenario in our experimental evaluation when evaluating speedups, our methodology is agnostic with respect to the integration choice, because we abstract architectural aspects when performing application mapping to analog/digital resource. From a technology perspective, both SRAM- and DRAM-based IMC architectures have been proposed [2]. Computation in these architectures can be achieved by taking advantage of charge sharing when simultaneously activating multiple memory rows. A promising alternative, which we focus on in this paper, is that of employing Non-Volatile Memory (NVM) devices, organized in a crossbar arrangement [2]. These devices, used to encode network weights in a differential mapping scheme, are employed at the junction of word- and bit-lines, and their resistances are modulated according to the target weight value at compile time.
As depicted in Fig. 2, AIMC crossbars encode inputs as word line voltages using Digital-to-Analog Converters, and currents, representative of MVMs results, are collected across bit lines into Analog-to-Digital Converters. Operations on AIMC crossbars are performed in parallel, and hence, can be realized in constant time complexity. For instance, a 256$\times$256 Phase Change Memory (PCM) crossbar is shown to perform a total of 65,536 MAC operations in 130ns  in Le Gallo et al.[12]. However, as computation is performed in the analog domain, outputs are affected by non-deterministic noise, deviating from expected results. This is due to a number of device and circuit non-idealities, including temporal and temperature-dependent conductance variations, device stochasticity, ADC mismatch, and IR drop [5].
Figure 2: (a) Depiction of an AIMC tile and its underlying compute mechanism. Each memristor, as depicted, is representative of a unit cell [^2] AIMC tiles may be realized using different architectures [13, 14], most prominently, pseudo-crossbar or conventional memory, where the input injection ways are different, and the single-bit-multi-bit encoding capabilities are also different. LionHeart is architecture agnostic..
Figure 3: Mapping of (a) FC and (b) CONV layers to device conductances. Weights are linearly scaled and mapped between $G_{min}$ and $G_{max}$. (c) Mapping/execution flow of a self-attention block when the maximum analog MAC ratio is achieved.
Figure 4: (a) High-level overview of our proposed framework. (b) The validation accuracy at the desired evaluation time ($t_{eval}=$1d) and training loss of a candidate AIMC layer during analog HWA retraining. Training is considered converged when, for a user-defined convergence window, the training loss does not decrease. (c) High-level overview of analog HWA retraining.
II-B Hardware-Aware Training
Analog HWA offline training is a common approach to mitigate the AIMC device and circuit non-idealities, by making the model more robust for deployment. When performing HWA training, weight-noise is injected during forward propagation passes. Mathematically, this is expressed as follows:
$y_{i}=\alpha^{\textrm{out}}_{i}f_{\textrm{adc}}\big{(}\sum_{j}({w}_{ij}+\sigma_{\textrm{w}}\xi_{ij})(f_{\textrm{dac}}(x_{j})+\sigma_{\textrm{inp}}\xi_{j})+\sigma_{\textrm{out}}\xi_{i}\big{)}+\beta_{i},$ (1)
where $f_{\text{adc}}$ and $f_{\text{dac}}$ model the analog-to-digital and digital-to-analog processes, with dynamic scaling and range clipping, and $\xi$ denotes Gaussian noise. $\sigma$ indicates noise sampled from a zero-centered normal distribution. In addition to training noise-resilient weights, it is also possible to train other parameters, such as the input (DAC) bounds of each tile, using back-propagation. We employ the open-source AIHWKIT toolkit [5] to expose these behaviours using system-level explorations in LionHeart, allowing both the realistic characterization of analog non-idealities and analog HWA retraining. While the toolkit, and by extension, the LionHeart framework can be used to model other device technologies, such as Resistive Random-Access Memory (RRAM), in this paper, we model PCM. The PCM device technology was chosen as compared to other device technologies, it utilizes fab–friendly materials and can achieve high conductance precision. Section IV-B3 details the specific assumptions made during analog HWA retraining.
II-C Computational Complexity and Execution of Convolutional and FC Layers Using IMC
FC layers execute a number of MAC operations which is linearly proportional to the number of input and output elements. Specifically, they have a complexity of $\mathcal{O}(MN)$, with $M$ and $N$ representing the number of input and output features, respectively. A CONV layer that applies $F$ filters to compute an output feature map of size $W_{o}\times{}H_{o}\times{}F$ requires a number of MAC operations in the order of $\mathcal{O}(W_{o}H_{o}CK^{2}F)$, where $K$ represents the kernel size and $C$ the number of input channels.
Weights of FC layers can be computed directly using the MVM operation, whereas weights of CONV layers can be transformed to MVMs using the im2col algorithm [15] (see Fig. 3). Using im2col, weights of CONV layers are unfolded and indexed as a matrix, where each column of the matrix contains the unrolled values of a filter. For both layer types, unfolded weight matrices are typically much larger than the attainable dimensions of AIMC tiles. For both layer types, unfolded weight matrices are typically much larger than the attainable dimensions of AIMC tiles, e.g., all convolutional layers of ResNet20 require 32 256$\times$256 tiles. Hence, tiling can be employed to partition large matrices into a number of discrete tiles. This strategy requires an additional computing step to aggregate partial results, which is performed per tile, rather than per element [4]. Unless otherwise stated, we utilize a distributed weight mapping scheme, in which larger matrix are evenly distributed among tiles.
For self-attention blocks, we consider mapping the first three FC layers of each head and the last FC layer after concatenation. FC layers within the scaled-dot product attention are not considered, as they have dynamic weights, which for AIMC require constant reprogramming during infererence.
II-D Related Work
Most related work, which performs application mapping on AIMC-accelerated systems, consider resource limitations as their main (or only) constraint. PUMA [16] employs compiler analysis passes to define which parts of the application have the most potential for speedup, while AERO [17] formulates this same problem as a cost function minimization. Potential issues introduced by non-idealities and noisy environments, characterizing AIMC devices, are overlooked in the above-mentioned works. They are similarly disregarded in [18] and [19], where all FC and CONV layers are executed using AIMC tiles, without attempting to control induced accuracy degradation. Some consideration of the effect of analog computing on accuracy is included in [7], where the authors propose a hybrid solution where the first and last layers of DNNs are executed on digital resources, postulating that these two layers are more sensitive to noise.
Two recent works present input channel-wise weight selection methods to map to heterogeneous analog-digital resources. The first, Harmonica [20], does so without employing any re-training strategy. The input is assumed to already be HWA trained, and some layers are confined to be implemented using digital resources. For channels of unconfined layers, the Hessian sensitivity [21] is determined. These are then ordered from the most sensitive to the least sensitive. The (next) most sensitive channel is converted to digital and the accuracy of the network is determined. This process is repeated until the accuracy of the network drops below a pre-defined threshold. Harmonica is described in full in Algorithm 1 of [20].
The second, ODiMO [22], relies on the unique approach of training a supernetwork, where for each channel, both digital and AIMC implementations are trained simultaneously. After the network is trained, for each channel, either the digital or AIMC implementation is pruned. Analog computation is modelled using 2-bit quantization. Both of these methods do not consider (i) the variability of analog tiles as well as (ii) the temporal evolution of analog weights (conductances).
The complexity of devising high-performance, but accuracy-constrained mappings, is compounded by the varying size and sensitivity towards noise of the different layers in a DNN structure. Moreover, the number of possible mappings makes exhaustive exploration impractical, even for simple cases, as discussed, in the context of hardware quantization, in [23]. For example, even when layer-wise mapping is considered, a relatively low-depth model like AlexNet, composed of just seven layers, would therefore present $2^{7}$ possible analog-digital combinations. All of these would have to be trained and tested to arrive at an exact solution to the mapping problem. In contrast to Harmonica and ODiMO, we employ a heuristic-based layer-wise mapping strategy, which scales linearly with respect to the number of layers, as it only considers one target candidate layer at a time.
III LionHeart framework
[…]
TABLE I: The number of layers, weights, and the top-1 (baseline, floating-point) accuracy of each evaluated network.
Model Weights Top-1 Acc./F1 (%)
(# Mappable Layers) CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
ResNet8 (10) 77,360 77,968 87.38 59.40
ResNet20 (20) 268,336 270,522 90.52 65.15
MobileNet (15) 3,195,328 3,206,282 90.15 63.05
VGG-16 (14) 14,715,584 14,761,664 92.79 69.51
AlexNet (7) 15,378,112 15,562,432 91.29 65.70
(# Mappable Layers) SQuAD SQuAD
MobileBERT (169) 24,844,5441 90.2
[…]
While not investigated in this paper, it is possible to extend the methodology of LionHeart to co-apply different techniques to further improve accuracy, e.g., tunable ADC resolutions in RAELLA [25].
IV Evaluation
IV-A Experimental Setup
For the considered DNNs, similarly to [7] and [8] we evaluate the performance of LionHeart on the CIFAR-10 dataset [26]. Moreover, we also provide outcomes on the more challenging CIFAR-100 dataset. [26]. We consider five DNN architectures targeting edge devices, as listed in Table I, where the size of the last FC layer is modified according to the number of output classes (10 and 100, respectively).
While the degree to which a DNN architecture is parameterized for a given dataset is difficult to determine quantitatively, the DNN architectures were selected such that they range from being under- to over-parameterized so that they are representative of a broader range of networks and chosen tasks. To demonstrate the generality of LionHeart beyond CNNs, we also apply it to the MobileBERT transformer benchmark, processing on the SQuAD dataset [27] [^3] We refer to the SQuAD 1.1 dataset as SQuAD throughout the paper..
We collect results using CIFAR-10 and CIFAR-100 for the first five DNN architectures targeting edge devices as listed in Table I. Additionally, we collect results using SQuAD for MobileBERT. The number of weights of these networks range from $\approx$78K to $\approx$25M. Data is split into training, validation, and testing sets. For CIFAR-10 and CIFAR-100, the validation dataset was obtained by splitting the original training dataset into two separate datasets, comprising 90% of training images and the remaining 10% of validation images. The same randomly sampled validation set was used for all experiments. For SQuAD, the pre-existing development set (comprising 10% of all inputs) was used as the validation set. The test sets of all datasets were solely used for final model evaluation.
Figure 5: For (a,b) ResNet8, (c,d) ResNet20, (e,f) AlexNet, (g,h) VGG16, (i,j) MobileNetV2, and (k,l) MobileBERT, the MAC ratio (as a percentage) and corresponding test set accuracies (F1 scores for SQuAD) at $t=$1d. The MAC ratio quantity represents the ratio between the number of MACs performed using AIMC and the total number of MACs
TABLE II: Shared training hyper-parameters for the CIFAR-10/ CIFAR-100 and SQuAD tasks.
Parameter CIFAR-10/CIFAR-100 SQuAD
Optimizer SGD + Momentum
Loss Function Cross Entropy
LR Scheduler CosineAnneling Linear
T_max 50 3
Batch Size 256 8
Digital
LR 0.057 0.05
Momentum 0.867 0.99
Weight Decay 0 0.00054
Analog
LR 0.024 0.02
Momentum 0.775 0.9
Weight Decay 0 0.003
IV-B Training and Evaluation Strategies
IV-B1 Initial hyper-parameter exploration
Our framework assumes that the input DNN is pre-trained in floating-point precision. The selected models differ in depth, size, and complexity, and hence require different hyperparameters to achieve optimal training performance. For all networks listed in Table I, we empirically determined two sets of hyper-parameters, pertaining to the analog and digital domain, respectively. Then, these were applied to the networks. These are listed in Table II. Further details on the adopted experimental setups in the two domains are provided in the following.
IV-B2 Digital training
For all Convolutional Neural Networks, networks were trained using floating-point precision until the training loss did not decrease for five consecutive epochs. For MobileBERT, a fixed number (3) of training epochs was used. All networks were trained with cross-entropy loss in conjunction with Stochastic Gradient Descent (SGD) with momentum.
IV-B3 Analog HWA re-training
To perform analog HWA re-training, the weights of the trained digital networks were employed as a starting point. The IBM AIHWKIT was used to inject noise during forward propagations. We refer the reader to [5] for a comprehensive tutorial on HWA training using IBM AIHWKIT. To configure the behavior of AIMC tiles and HWA training, the IBM AIHWKIT utilizes different configurations which are defined using a standardized data class data structure. A number of pre-defined configurations are provided. We used the provided InferenceRPUConfig() configuration and PCMLikeNoiseModel phenomenological inference model.
The following additional modifications were made to the simulation configuration: (i) biases are assumed to be digital (i.e., they are not encoded on the last column of AIMC tiles). (ii) Channel- (i.e., column-) wise weight scaling, which has a negligible performance impact, is performed, where weights are mapped to a per-channel conductance range during learning after each mini-batch. (iii) The size of each AIMC tile is assumed to be 256x256. (iv) Layers that span across multiple AIMC tiles are assumed to be split evenly. (v) During training, multiplicative Gaussian noise is applied to unit weight values, with a standard deviation of $0.08$, extracted from hardware measurements provided by the authors of [12].
IV-B4 Accuracy evaluation
As AIMC hardware is inherently stochastic, multiple evaluation instances are required to determine representative performance metrics. Consequently, accuracy was measured 20 times and averaged across experiments to obtain the values reported in Section IV-D. We considered five user-defined accuracy [^4] For SQuAD, instead of considering the accuracy, the F1 score is considered as a percentage value. thresholds of [0.5%, 1.0%, 3.0%, 5.0%, 10.0%]. Moreover, analog NVM devices such as PCM are susceptible to temporal conductance drift. Since the behavior of AIMC hardware can evolve over time, we explore in Section VI accuracy metrics for AIMC accelerators programmed considering drift levels predicted at [one second (i.e., $t=1s$), one minute (i.e., $t=60s$), one hour (i.e., $t=3600s$), one day (i.e., $t=86400s$)]. These are then evaluated at different points in time, to assess the effect of a mismatch between predicted and actual drift.
IV-B5 Full System Evaluation
The target simulated system for this work is a high-performance edge processor with an in-order ARM CPU core running at a frequency of 2.3GHz, with a 64KB each L1 data and instruction caches, 1MB of L2 cache, and 8GB of 2400MHz DDR4 RAM, as detailed in Table IV. The software stack for the target DNNs includes the DNNs accelerated via the Eigen C++ library, running in user-space atop the Linux 5.4 kernel, and an Ubuntu 16.04 LTS disk image. Each implementation resulting from a LionHeart mapping is run for 10 inferences to reduce simulation noise. Measured execution time deviations among different runs of the same mappings never exceeded 4% of the run-time.
IV-C Baselines
LionHeart (in all graphs, referred to as LH) is evaluated by comparing the accuracy and performance achieved with respect to two baselines: (i) Fully digital. The floating-point DNN model input of our methodology, where none of its layers is accelerated using AIMC. It achieves the highest accuracy, but presents the longest run-time. (ii) Fully-analog. A DNN implementation executing all FC and CONV layers with static weights using AIMC. This baseline represents the opposite of the fully-digital: it maximizes performance, while potentially resulting in very large accuracy degradation. In Section V, we provide a further comparative evaluation with state of the art strategies, namely Ueyoshi et al. [7] (i.e., DIANA) and Harmonica [8].
IV-D Accuracy Evaluation
For five different drop thresholds and baseline configurations, we determined the ratio between MACs performed using AIMC and the total number of MACs for CIFAR-10, CIFAR-100 and SQuAD (see Fig. 5). As expected, for all networks the All Digital configuration has the highest test set accuracy and the All Analog configuration has the lowest test set accuracy. The proposed LionHeart configurations exhibit a gradual transition from the accuracy of the largest to the smallest test set and a corresponding decrease in the digital/analog MAC ratio between all configurations. Thus, it enables exploration of the performance/accuracy space exposed by AIMC acceleration. Critically, the design space is navigated in a controlled way, that is, abiding to a user-defined accuracy degradation constraint. In this way, configurations exhibiting both high performance (high ratio of AIMC acceleration) and high accuracy can be derived. As an example, almost 60% of the computation can be AIMC-accelerated in the ResNet-8 network for an accuracy threshold of 5%, while an all-analog alternative would have a decrease in accuracy of 20%.
For a small number of networks and datasets, it is observed that the All Analog configuration has a higher test set accuracy than some of the LionHeart configurations; particularly those with larger drop thresholds, e.g., 10%. Moreover, some LionHeart configurations with larger drop thresholds have a higher accuracy than others with smaller drop thresholds. To this end, we make the following observations. First, heuristic-based methods are inherently stochastic in outcome. While we investigate the variability of LionHeart and related work in Section V, in this section, we first investigate the one-shot behavior of LionHeart in earnest. Second, LionHeart enforces a global drop threshold, meaning that if a network fails to converge when a layer is retrained using HWA training, a large accuracy drop can be incurred (as large as this threshold). This can result in a sub-optimal outcome if one of the first layers incurs a large accuracy drop, as all subsequent layers are subjected to the remaining drop threshold.
Lastly, given a fixed accuracy threshold, LionHeart does not penalize configurations for being close to this threshold. This means that while in practice it would not make sense to specify, if an accuracy drop of Xwas indeed specified, and the accuracy drop of the fully-analog configuration was X-N%, the resulting mapping would not be penalized for being between X-Nand X%.
TABLE III: Ratio of configurations where a layer is AIMC-accelerated, per layer type and topological location in the MobileBERT encoder.
Encoder Block Location First Middle (Avg.) Last
Encoder Layer Average Retention in Analog (%)
Bottleneck 45 52 46
Attention 78 81 61
FFN0 80 68 56
FFN1 75 74 62
FFN2 68 70 66
Intermediate 48 75 48
Output 45 52 40
We observe that no obvious pattern emerges regarding the topological order of layers in our explorations, other than that CONV layers are more likely to be AIMC-accelerated compared to FC ones. For the same layer type, whether or not a layer will be executed in the analog domain depends instead on the trained weights, and not on the layers and the network topology, thus mandating exploration heuristics such as LionHeart. A similar conclusion can be reached for the MobileBERT case. Again, the choice of accelerated layers does not follow a clear topological pattern, as we report in Table III, where the conversion percentage is reported for the first and last, in addition to the average of all middle encoder blocks. The first attention layers are more amenable to be executed in AIMCs, but this is not the case for Bottleneck layers.
From our analysis, for both CNNs and transformer-based networks, we observe that the first layer is usually one of the most sensitive. The intermediate layers are generally seen to be more sensitive than the last layer. Our analyses demonstrate that, especially for larger networks with a more complex topology, properly mapping layers to digital or analog resources is non-trivial.
TABLE IV: Architectural parameters for full system simulation.
CPU Model 1x In-Order @ 2.3GHz
ISA ARMv8
Supply Voltage VDD 1.3V
L1 Data/Instruction Cache 64kB
L2 Cache 1MB
Memory 8GB DDR4 @ 2400MHz
AIMC Tile Latency 100ns
AIMC Tile Data Transfer Bandwidth 4GB/s
AIMC Tile Energy Efficiency 20 TOp/s/W
CPU + SIMD (Baseline) Peak Efficiency 22.8 GOp/s/W
Figure 6: (a) Speedup and (b) energy efficiency gain resulting from different application mapping strategies, when performing an inference on benchmark DNNs. Data is normalized with respect to a fully digital execution with the ARM NEON SIMD co-processor in lieu of AIMC acceleration.
IV-E Performance Evaluation
In LionHeart, the workload associated with DL models and their layers are assumed to be architecture-agnostic, i.e., the ratio of analog MAC to total MAC operations is used as a proxy for speedup and energy efficiency.
This approach allows to derive hybrid mappings in a hardware-agnostic way, while also avoiding impractically lengthy simulations. On the other hand, it does not allow to directly quantify the performance and efficiency of mappings on a target architecture. To inspect this aspect, and also validate the link between MAC ratio and performance/efficiency for one system configuration, we herein execute entire inferences for the three largest CNN benchmarks (AlexNet, VGG16 and MobileNetV2) and measure the ensuing run-times, including computations not pertaining to FC nor CONV layers. We investigate mappings derived by LionHeart, with varying maximum accuracy degradation thresholds.
As a test vehicle, we adopted a simulated system featuring a single-core in-order CPU clocked at 2.3GHz and a two-level cache hierarchy, modelled within the ALPINE framework [4]. The architecture embeds tightly-coupled AIMCs, governed by dedicated extensions to the ARM instruction set supported by the CPU. Table IV lists the architectural parameters of the test system, which mirror those considered in [4] for the single-core, higher performance scenario.
To align the performance experiments with state-of-the-art low-power edge-domain devices, all DNN implementations are quantized to INT8 precision. The fully digital baseline architecture utilizes the ARM Neon SIMD co-processor in lieu of the tightly-coupled AIMC tiles. Note that due to the effects of memory transactions and latencies, the effective Op/s of both the AIMC-enabled and baseline systems are a fraction of the peak Op/s, at 57% and 35%, respectively.
The CPU in the evaluated system architecture uses customized ISA extensions to access AIMC tiles that are tightly coupled to the CPU cores, without requiring the traversal of the memory hierarchy for data transmission [4]. The extension instructions are implemented using unused opcodes in the ARMv8 architecture. During inference, three instructions are used: CM_QUEUE, which places packed inputs in the input memory of an selected AIMC tile, CM_PROCESS, which operates the AIMC tiles and performs MVMs, and CM_DEQUEUE, which retrieves the outputs from selected AIMC tiles and places them in destination registers. For all speedup and efficiency gain factor evaluations, it is assumed that a single CPU core is used and the required number of tiles to map all analog weights for the evaluated network are tightly coupled to this core.
In Fig. 6, we report the obtained speedups (with respect to fully-digital baselines) and efficiency gain factors. In order to align the performance experiments with state-of-the-art low-power edge-domain devices, all DNN implementations are quantized to int8 precision. Additionally, the fully digital baseline architecture utilizes the ARM Neon SIMD co-processor in lieu of the tightly-coupled AIMC tiles. Note that due to the effects of memory transactions and latencies, the effective Op/s of both the AIMC-enabled and baseline systems are a fraction of the peak Op/s, at 57% and 35% respectively.
These results highlight that the achievable speed-ups are highly dependent on the DNN structure. In particular, even when all layers are computed in the analog domain, which is an upper bound for speedup and efficiency in our setting, only a speedup of 24% can be reached in the MobileNetV2 benchmark. Such (comparatively) low run-time performance improvement is caused by the lower degree of data reuse of the depthwise separable convolutions employed in this network, which shift the balance between computation and memory access requirements toward the latter, limiting the benefit of accelerating MACs. However, even in this case, LionHeart, with a limited maximum accuracy degradation of 5%, still captures almost all of the available speedup and energy efficiency gains.
Figure 7: Approximate Pareto front of the achieved LionHeart mappings, and operational points of both variants of the Harmonica and FLMS methods. For each network, the elbow for LionHeart, i.e., the optimal operating point, is highlighted using a darker hue value.
AlexNet and VGG16, which do not employ separable convolutions, instead present more potential speedup (660% in the full-analog case for VGG16, 550% for AlexNet), due to the higher locality of the employed computations (which leads to less cache misses and memory-induced stalls). Even in those cases, LionHeart can harness a large part of the potential speedup if enough accuracy leeway is allowed. For a 5% accuracy degradation, a 425% speedup was measured in AlexNet and a 285% one in VGG16. AlexNet proved to be the most robust of the investigated networks. Finally, it can be noticed that system-level energy gains closely follow speedups. Indeed, AIMCs tiles are not a major contributor to overall system power (as opposed to processors and memories). Instead, they shorten run-time, ultimately lowering energy requirements. This is achieved both by allowing the execution of MVMs in constant time and by encoding DNN weights in the crossbar structure itself, hence easing the pressure on the memory hierarchy.
Figure 8: For the optimal LionHeart configuration, and both variations of the Harmonica and FLMS methods, (a-e) the MAC ratio (as a percentage) between the number of MACs performed using AIMC and the total number of MACs, and (f-j) the corresponding test set accuracies at $t=$1d. All experiments are repeated $n=5$ times.
TABLE V: Optimized hyper-parameters used for DIANA and Harmonica.
Parameter ResNet8 ResNet20 MobileNet VGG-16 AlexNet
Optimizer SGD + Momentum
Loss Function Cross Entropy
LR Scheduler CosineAnneling
T_max 200
Batch Size 64
Digital
LR 0.0945 0.00184 0.01866 0.0516 0.0157
Momentum 0.86 0.99 0.94 0.88 0.9
Weight Decay 0.00045 0.00017 0.00085 0.00058 0.00011
Analog
LR 0.0617 0.0882 0.0898 0.00619 0.00529
Momentum 0.9 0.82 0.9 0.9 0.99
Weight Decay 0.0008 0.00042 0.00037 0.00037 0.00091
V Comparison to Related Work
We compare against two established heuristic-based methods, DIANA [7] and Harmonica [8]. Comparisons are made using the CIFAR-10 dataset, as this dataset was originally used for evaluation by both of these methods. For DIANA (referred to as First-Last Mapping Strategy (FLMS) to encompass the more generic approach of executing only the first and last layer in high-precision), we consider two scenarios: (i) where the original network is in floating-point precision, akin to LionHeart, and (ii) where the original network is HWA-trained.
For Harmonica, we only consider layer-wise mappings, as channel-wise mappings are too computationally expensive, i.e., they require exploration of a much larger search space dependent on the number of channels rather than the number of layers, to be explored exhaustively. Two scenarios are also considered: (i) where the mapped network is not retrained (Harmonica) and (ii) where the mapped network is retrained using HWA retraining (Harmonica_T). The input to Harmonica is assumed to be HWA-trained. This is problematic, as HWA training requires exhaustive hyper-parameter optimization to achieve optimal results [5] when all layers are implemented using AIMC hardware. This is time demanding and computationally expensive. We avoid comparisons with ODiMO due to their use of a channel-wise split, which markedly reduces the analog utilization, as evidenced by their lower analog mac ratio [22].
To establish strong baselines for downstream comparison to LionHeart, for each DNN architecture, we performed hyper-parameter optimization for HWA training using the Optuna [28] library. Specifically, we employed the Tree-structured Parzen Estimator (TPESampler) and the Median pruning algorithm (MedianPruner) optimization algorithms over 10 trials. The resulting new hyper-parameters are listed in Table V.
FLMS and Harmonica do not determine mappings for different MAC ratio and accuracy values. Consequently, we determine the optimal LionHeart mappings by generating an approximate Pareto front in Fig. 7. Additionally, we determine and compare the average operational point for each method (FLMS and Harmonica). It is observed that the operational points of these methods are captured either on or underneath LionHeat’s Pareto front, meaning that LionHeart discovers the same or even better mapping configurations along its front.
To perform further comparisons using a single operational point in Fig. 8 for each method including LionHeart, we select the elbow for each network. The elbow method transverses operational points along the Pareto front from left-most point (where the quantity encoded using the y-axis is maximized) to the right-most point (where the quantity encoded using the x-axis is maximized). The optimal operation point, or elbow, is distinguished by the fact that before reaching it, the y-axis quantity remains almost unchanged, and after reaching it, rapidly decreases [29].
Figure 9: The impact of temporal drift when varying $t_{eval}$ in [0, 86400] for the ResNet8 network evaluated on CIFAR-100.
In Fig. 8, it can be observed that for all networks, both variants of FLMS result in a high (near 100%) MAC ratio with zero variance. Although the resulting accuracy values have a small variance, they are unpredictable, i.e., compared to other configurations, the average test accuracy varies greatly across networks. Moreover, for most (4/5) networks, compared to other configurations, the accuracy values are the lowest.
On average, both variants of Harmonica yield higher accuracy values when compared to FLMS, however, the average MAC ratios are low and the variability of both the MAC ratios and accuracy values are larger. When the mapped networks from Harmonica are retrained using HWA retraining, i.e., for Harmonica_T, the average accuracy values increase. It is noted that these networks exhibit similarly variability to the networks originally generated from Harmonica.
LionHeart consistently achieves competitive accuracy values with reduced variation compared to both Harmonica variants, for all networks expect ResNet20, while maintaining relatively high MAC ratios. For the single network (ResNet20) where LionHeart performs sub-optimally, it has a significantly higher MAC ratio with respect to other configurations (not including FLMS). It is speculated that the reason Harmonica does not perform at least as well with respect to accuracy, is that at the beginning, in contrast to LionHeart, it requires an already HWA retrained network, and does not retrain any layers. LionHeart begins with a stronger baseline (a pre-trained floating-point network) and performs retraining, providing an opportunity for accuracy losses to be recovered when each layer is converted to analog.
From Fig. 7, if another point along the Pareto front is taken with a smaller MAC ratio, the resulting test precision significantly improves, that is, for a MAC ratio of 85.63%, the test accuracy is 88.36%. This is larger than the average test accuracy for the other aforementioned configurations.
VI Temporal Drift and Evaluation Time Analysis
We study the impact of varying the evaluation time, $t_{eval}$, used during analog HWA retraining on the simulated inference time of the system. In Fig. 9, these metrics are reported for all configurations (except All Digital, which does not have a temporal dimension) using the ResNet8 network and CIFAR-100 dataset. For both sets of mappings, mostly unique intermediate layers are retained in analog. Due to the (relatively) small drop threshold, the first and last layers for each are in digital. It can be seen that the test accuracy decreases as a function of the drift time. Moreover, as the proportion of analog MACs increases, the robustness to drift decreases. Most notably, for the All Analog configuration, independent of the evaluation time, the test accuracy decreases significantly ($>$10%). For all other configurations, no significant degradation is observed until $t>$1h. However, the temporal dependence of the test set accuracy is effected. Unsurprisingly, when a larger evaluation time is specified, the initial test set accuracy (at $t$=0) is also larger. While varied, for most configurations the robustness to drift is observed to improve as the evaluation time is increased. This is significant because, when our methodology is used, a hard constraint on the average behavior of the drop in accuracy can be enforced at a given drift (evaluation) time.
VII Comparison to Standard Hardware-Aware Retraining
Figure 10: Approximate Pareto front of the achieved LionHeart mappings, operational points of both variants of the Harmonica and FLMS methods, and the analog hardware-aware retraining ceiling for ResNet8 evaluated on CIFAR-10.
Finally, to compare LionHeart and other methods to standard hardware-aware retraining, we retrain all 1,024 mappings of ResNet8 and compute the exact Pareto front in Fig. 10, which is indicative of the ceiling or best case scenario when all analog layers are retrained at the same time using the same convergence criteria for all other experiments. It can be seen that for MAC ratio values smaller than 80%, LionHeart is extremely effective. For a MAC ratio of $\approx$60%, LionHeart exceeds this ceiling, indicating that progressively training analog layers can result in increased accuracy. This is in agreement with prior work [30].
VIII Conclusion
In this paper, we have presented LionHeart, an accuracy-driven framework for mapping DL workloads on hybrid analog-digital systems. The effectiveness of our method was demonstrated by deploying five different DNNs trained for image classification to heterogeneous Processing Units. Both the MAC ratio between analog and digital operations and the effect of temporal drift were considered to enforce varying user-specified constraints. Performance evaluation was carried out for three DNNs using a simulated ARMv8-based processor with tightly coupled AIMC tiles. The optimal mappings of LionHeart were compared to related work, and it was demonstrated that LionHeart achieved state-of-the-art results. Finally, additional analysis was performed by investigating the effects of temporal drift and the variation of the evaluation time. Future work entails the hardware-based architecture-level exploration and investigation of the additional training time/complexity overhead that is incurred.
Acknowledgments
This work was supported in part by the the Swiss NSF Edge-Companions project (GA No. 10002812); in part by the EC H2020 FVLLMONTI Project under Grant 101016776; in part by the ACCESS—AI Chip Center for Emerging Smart Systems, sponsored by InnoHK funding, Hong Kong, SAR; in part by the Swiss State Secretariat for Education, Research, and Innovation (SERI) through the SwissChips Research Project; and in part by EC H2020 WiPLASH under Grant 863337.